# Forbal: Force Balanced 2-5 Degree of Freedom Robot Manipulator Built from a Five Bar Linkage

### v1

@@ II-A Force balance of a planar five bar linkage
The five bar linkage is modeled as made of two $R$ chains that connect at $\bm{p}_{c}$. Joints and links are identified by a subscript ijij, where $i$ denotes the chain number, and $j$ denotes the joint and connected link in that chain, starting from the base. The first joint in each of the two chains is aligned with the $z$-axis of the fixed frame and placed at a height of $l_{h}$ on the base.
@@
where $m_{ij}$ and $\dot{\bm{r}}_{ij}$ are the masses and the Cartesian linear velocity vector of the CoM of link ijij about the fixed frame, respectively.
@@ II-B Forbal-2: Planar 2-DOF manipulator
Starting from the model of Fig. 3, a force balanced five bar linkage (Forbal-2) was developed. All the links are of the same length $l$, and the first joints of the two $R$ chains are placed on the same axis ($l_{d}=0$). All links have the same length $l$ between joints which simplifies the kinematics, maximizes workspace, and ensures that the manipulator does not reach singular configurations within joint limits. Counter masses are only required on 3 links to achieve force balancing for the 5BL, and the total mass is the same or lower compared to placing them on all 4 links according to our previous analysis [15]. We attach them on Links 11, 21, and 22 through extending the profile behind the direction of the kinematic chain .
@@ II-C Forbal-5: Extension to Spatial 5-DOF manipulator
The static end-effector payload of Forbal-2 is replaced with a 2 DOF $E$ motor. These two additional rotational joints (named Joint 3 and Joint 4) do not intersect each other due to space constraints. Joint 3 is placed on the Forbal-2 end effector position $\bm{p}_{e}$, and its direction is parallel with Joints 11 and 21. Joint 4 is orthogonal to it, that actuates the $E$ implement.
The $E$ motor and implement together constitute a new end effector payload which should have an inline center of mass, aligned with Link 22 for full force balance, assuming there are no frictional forces acting on the system from the joints and actuator gears.
@@ III-A Planar Geometric Controller
Given a fixed frame position vector for the end-effector in the xzxz plane $\bm{p}_{e}\in\mathbb{R}^{2}$, the inverse kinematics derive the required corresponding joint angle configuration $q_{11}$ and $q_{21}$. First, we calculate the vector from the position of the servomotors in the xzxz plane, $\bm{p}_{b}$ to the end-effector: $\bm{p}_{be}=\bm{p}_{e}-\bm{p}_{b}$. We then derive the end-effector Euclidean distance $\bm{p}_{be}||=\sqrt{p_{be_{x}}^{2}+p_{be_{z}}^{2}}$ and its angle with respect to the $x$ axis at $\bm{p}_{b}$ as $\theta_{be}=\tan^{-1}(\frac{p_{be_{z}}}{p_{be_{x}}})$.
@@ III-B Forbal-5: Extension of planar IK to 5-DOF pose
With the addition of 2-DOF motion, the end-effector implement can be set to a desired 5-DOF pose. We define the $E$ implement pose as $\bm{T}_{e}=\begin{bmatrix}{}_{O}p_{e_{x}}&{}_{O}p_{e_{y}}&{}_{O}p_{e_{z}}&{}_{B}p_{\beta}&{}_{O}p_{\gamma}\end{bmatrix}^{\intercal}$. The first three components are the Cartesian coordinates in the global frame $_{O}\bm{p}_{e}$. The next two orient the end-effector; $_{B}p_{\beta}$ is the pitch angle of the $E$ motor about the 5BL plane $y_{B}$ axis orthogonal to $x_{B}z_{B}$ at $_{O}\bm{p}_{m}$ (Joint 3 position). The yaw angle $_{O}p_{\gamma}$ is about the global z-axis after the pitch rotation is applied. The final $E$ implement orientation in the global frame is $R_{y}(_{B}p_{\beta})R_{z}(_{O}p_{\gamma})$. The diagram illustrating the kinematics of Forbal-5 is shown in Fig. 7.
Figure 7: Forbal-5 manipulator. The fixed origin axes are noted as $x_{O}$, $y_{O}$ and $z_{O}$, and the Joint 0 rotated 5BL base frame is shown with the $x_{B}z_{B}$ plane. The end effector position in world frame is $_{O}\bm{p}_{e}$, and $_{B}\bm{p}_{bm}$ is the vector from the center of Joints 11 and 21 to the $E$ motor in the $x_{B}z_{B}$ plane.
@@
Since the origins of the reference frames of the two motors are along the axis $z_{0}$,and the center of the $E$ motor is fixed to the center of the end effector, Joint 4 is directly positioned on the 5BL’s $x_{B}z_{B}$ plane, allowing for a simple inverse kinematics formulation that utilizes the previous planar Forbal-2 inverse kinematics with some modifications.
@@
Note that the y-axis position of the $E$ implement always lies on the $x_{B}z_{B}$ plane, which rotates about the global z-axis $z_{O}$.
We have to adjust the position of the implement about the pitch angle $p_{\beta}$ to find the desired coordinates of the $E$ motor in $x_{B}z_{B}$:
@@
where $_{B}R^{-1}(_{B}p_{\beta})\in SO(3)$ is the Y-axis rotation caused by the desired pitch $_{B}p_{\beta}$, and $_{M}\bm{p}_{me}$ is the vector from Joint 3 to Joint 4 in the $E$ motor frame.
To use the Forbal-2 inverse kinematics, we define the target position of the $E$ motor from the center point of Joint 11 and Joint 21 (the mount axis) in the 5BL frame $_{B}\bm{p}_{bm}={{}_{B}\bm{p}_{Om}}-\bm{p}_{b}$. Using this formulation, (11)-(12) can be used to find $q_{11}$ and $q_{21}$. The $E$ motor joint angle $q_{3}$ simply adjusts for the desired pitch $p_{\beta}$ about $y_{B}$ from $x_{B}$ according to the calculated inverse joint angles: $q_{3}=q_{22}-q_{21}+{}_{B}p_{\beta}$.
@@ IV-A Prototype
$E$ (Forbal-2) 43.7 10.99 0.85 0.00
$E$ (Forbal-5) 27.1 2.44 1.38 0.00
$E$ motor 102.0 12.0 4.90 0.01
$E$ implement 3.02 0.00 0.00 6.37
@@
For Forbal-5, we selected the high torque Dynamixel XM430-W350-T motor for Joint 0, and a 2-axis Dynamixel 2XL430-W250-T motor to implement Joints 3 and 4 in the $E$ motor. Due to the placement of joints in the motor, once Joint 3 rotates the 5BL is no longer inline and theoretically the masses are only partially balancing the payload. Here we found again that the friction from the passive joints and gears in the motor facilitate force balancing.
@@ IV-F Discussion of Forbal-5 experimental results
The major difference between Forbal-5 and Forbal-2 from a force balancing perspective is an increase in the end-effector payload. While the mechanical design of the $E$ motor with the additional 2 joints is designed to be as close to inline for which the counter mass placement is valid, in theory it is still not possible to perfectly force balance the design with the fixed counter mass positions of the base Forbal-2 design.

### v2

@@ II-A Force balance of a planar five bar linkage
The five bar linkage is modeled as made of two RR chains that connect at $\bm{p}_{c}$. Joints and links are identified by a subscript $ij$, where $i$ denotes the chain number, and $j$ denotes the joint and connected link in that chain, starting from the base. The first joint in each of the two chains is aligned with the $z$-axis of the fixed frame and placed at a height of $l_{h}$ on the base.
@@
where $m_{ij}$ and $\dot{\bm{r}}_{ij}$ are the masses and the Cartesian linear velocity vector of the CoM of link $ij$ about the fixed frame, respectively.
@@ II-B Forbal-2: Planar 2-DOF manipulator
Starting from the model of Fig. 3, a force balanced five bar linkage (Forbal-2) was developed. All the links are of the same length $l$, and the first joints of the two RR chains are placed on the same axis ($l_{d}=0$). All links have the same length $l$ between joints which simplifies the kinematics, maximizes workspace, and ensures that the manipulator does not reach singular configurations within joint limits. Counter masses are only required on 3 links to achieve force balancing for the 5BL, and the total mass is the same or lower compared to placing them on all 4 links according to our previous analysis [15]. We attach them on Links 11, 21, and 22 through extending the profile behind the direction of the kinematic chain .
@@ II-C Forbal-5: Extension to Spatial 5-DOF manipulator
The static end-effector payload of Forbal-2 is replaced with a 2 DOF EE motor. These two additional rotational joints (named Joint 3 and Joint 4) do not intersect each other due to space constraints. Joint 3 is placed on the Forbal-2 end effector position $\bm{p}_{e}$, and its direction is parallel with Joints 11 and 21. Joint 4 is orthogonal to it, that actuates the EE implement.
The EE motor and implement together constitute a new end effector payload which should have an inline center of mass, aligned with Link 22 for full force balance, assuming there are no frictional forces acting on the system from the joints and actuator gears.
@@ III-A Planar Geometric Controller
Given a fixed frame position vector for the end-effector in the $xz$ plane $\bm{p}_{e}\in\mathbb{R}^{2}$, the inverse kinematics derive the required corresponding joint angle configuration $q_{11}$ and $q_{21}$. First, we calculate the vector from the position of the servomotors in the $xz$ plane, $\bm{p}_{b}$ to the end-effector: $\bm{p}_{be}=\bm{p}_{e}-\bm{p}_{b}$. We then derive the end-effector Euclidean distance $\bm{p}_{be}||=\sqrt{p_{be_{x}}^{2}+p_{be_{z}}^{2}}$ and its angle with respect to the $x$ axis at $\bm{p}_{b}$ as $\theta_{be}=\tan^{-1}(\frac{p_{be_{z}}}{p_{be_{x}}})$.
@@ III-B Forbal-5: Extension of planar IK to 5-DOF pose
With the addition of 2-DOF motion, the end-effector implement can be set to a desired 5-DOF pose. We define the EE implement pose as $\bm{T}_{e}=\begin{bmatrix}{}_{O}p_{e_{x}}&{}_{O}p_{e_{y}}&{}_{O}p_{e_{z}}&{}_{B}p_{\beta}&{}_{O}p_{\gamma}\end{bmatrix}^{\intercal}$. The first three components are the Cartesian coordinates in the global frame $_{O}\bm{p}_{e}$. The next two orient the end-effector; $_{B}p_{\beta}$ is the pitch angle of the EE motor about the 5BL plane $y_{B}$ axis orthogonal to $x_{B}z_{B}$ at $_{O}\bm{p}_{m}$ (Joint 3 position). The yaw angle $_{O}p_{\gamma}$ is about the global z-axis after the pitch rotation is applied. The final EE implement orientation in the global frame is $R_{y}(_{B}p_{\beta})R_{z}(_{O}p_{\gamma})$. The diagram illustrating the kinematics of Forbal-5 is shown in Fig. 7.
Figure 7: Forbal-5 manipulator. The fixed origin axes are noted as $x_{O}$, $y_{O}$ and $z_{O}$, and the Joint 0 rotated 5BL base frame is shown with the $x_{B}z_{B}$ plane. The end effector position in world frame is $_{O}\bm{p}_{e}$, and $_{B}\bm{p}_{bm}$ is the vector from the center of Joints 11 and 21 to the EE motor in the $x_{B}z_{B}$ plane.
@@
Since the origins of the reference frames of the two motors are along the axis $z_{0}$,and the center of the EE motor is fixed to the center of the end effector, Joint 4 is directly positioned on the 5BL’s $x_{B}z_{B}$ plane, allowing for a simple inverse kinematics formulation that utilizes the previous planar Forbal-2 inverse kinematics with some modifications.
@@
Note that the y-axis position of the EE implement always lies on the $x_{B}z_{B}$ plane, which rotates about the global z-axis $z_{O}$.
We have to adjust the position of the implement about the pitch angle $p_{\beta}$ to find the desired coordinates of the EE motor in $x_{B}z_{B}$:
@@
where $_{B}R^{-1}(_{B}p_{\beta})\in SO(3)$ is the Y-axis rotation caused by the desired pitch $_{B}p_{\beta}$, and $_{M}\bm{p}_{me}$ is the vector from Joint 3 to Joint 4 in the EE motor frame.
To use the Forbal-2 inverse kinematics, we define the target position of the EE motor from the center point of Joint 11 and Joint 21 (the mount axis) in the 5BL frame $_{B}\bm{p}_{bm}={{}_{B}\bm{p}_{Om}}-\bm{p}_{b}$. Using this formulation, (11)-(12) can be used to find $q_{11}$ and $q_{21}$. The EE motor joint angle $q_{3}$ simply adjusts for the desired pitch $p_{\beta}$ about $y_{B}$ from $x_{B}$ according to the calculated inverse joint angles: $q_{3}=q_{22}-q_{21}+{}_{B}p_{\beta}$.
@@ IV-A Prototype
EE (Forbal-2) 43.7 10.99 0.85 0.00
EE (Forbal-5) 27.1 2.44 1.38 0.00
EE motor 102.0 12.0 4.90 0.01
EE implement 3.02 0.00 0.00 6.37
@@
For Forbal-5, we selected the high torque Dynamixel XM430-W350-T motor for Joint 0, and a 2-axis Dynamixel 2XL430-W250-T motor to implement Joints 3 and 4 in the EE motor. Due to the placement of joints in the motor, once Joint 3 rotates the 5BL is no longer inline and theoretically the masses are only partially balancing the payload. Here we found again that the friction from the passive joints and gears in the motor facilitate force balancing.
@@ IV-F Discussion of Forbal-5 experimental results
The major difference between Forbal-5 and Forbal-2 from a force balancing perspective is an increase in the end-effector payload. While the mechanical design of the EE motor with the additional 2 joints is designed to be as close to inline for which the counter mass placement is valid, in theory it is still not possible to perfectly force balance the design with the fixed counter mass positions of the base Forbal-2 design.